# Towards Quantum Universal Hypothesis Testing thanks: B. Chen would like to acknowledge the support of a Google gift fund

### v1

@@ III-C Qudit Independent Measurements
Unfortunately, extending the above result to higher-dimensional Hilbert spaces gives an error exponent that scales poorly with the dimension of the space. For example, any Hermitian operator on $b$ qubits ($d=2^{b}$) can be decomposed using $d^{2}-1$ components, the Pauli strings [39]. Then applying the same method would give
@@
However, for general $d$-dimensional qudit states, we can do better using independent measurements by applying the following result from quantum state tomography [40]:
@@
$\mathbb{P}\left[\left\|\tau-\hat{\tau}\right\|_{1}\geq\delta\right]\leq d\exp\left\{-\frac{m\delta^{2}}{86dr^{2}}\right\}.$ (23)
Note that this is a special case of Guţă’s [40] more general result, where they take measurements that are $2$-designs [40, Definition 1]. Additionally, since $r\leq d$, we can weaken the bound to the following:
$\mathbb{P}\left[\left\|\tau-\hat{\tau}\right\|_{1}\geq\delta\right]\leq d\exp\left\{-\frac{m\delta^{2}}{86\cdot d^{3}}\right\}.$ (24)
This yields a concentration result that we can use in the same manner as before.
@@
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86d^{3}}\right\},$ (26)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\log(1/\alpha)\right)$ term in the exponent.
For proof see Section A-F. This also yields a result for the two-sample QUHT setting.
@@
$\beta_{\rho,\sigma}(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344d^{3}}\right\},$ (28)
where $\lesssim$ is hiding an additive $O(k^{1/2}\ln(1/\alpha))$ term in the exponent.
See Section A-G for proof.
@@ III-D Entangled Measurements
While independent measurements are easier to implement, it has been shown that using entangled measurements allows for quantum state tomography with a sample complexity that is unachievable using independent measurements [41].
@@
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\},$ (32)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\log^{1/2}(m)\right)$ term in the exponent.
See Section A-H for the proof. We can also use this to get results for two-sample universal hypothesis testing. In particular, using entangled measurements, we can achieve the following:
@@
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\},$ (34)
where $\lesssim$ hides an additive $O(k^{1/2}\ln^{2}(k))$ term in the exponent.
See Section A-I for the proof.
@@ IV Discussion
Recall that in the classical universal setting, Hoeffding was able to characterize the error exponent in terms of the KLD. In this work, we have established positive error exponents in terms of not the KLD but the trace norm. Specifically, we showed various results of the following form:
$\lim_{n\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(M_{m}))\right\}=-m\frac{\left\|\rho-\sigma\right\|_{1}^{2}}{c},$ (35)
@@
Comparing with the classical results for the universal hypothesis testing problem, which achieves an error exponent scaling as the relative entropy between $\rho$ and $\sigma$, this result appears somewhat lacking. For example, we have quantum Pinsker’s inequality [42, Theorem 3.1], which establishes that the squared trace distance forms a lower bound for the relative entropy. Since Hoeffding establishes his result in the classical setting using a large deviation type bound, a first attempt at a quantum generalization might utilize quantum variations of the classical Sanov’s theorem [6, Theorem 11.4.1]. However, in [14] and [16], the result required knowledge of the alternative hypothesis, which is not reasonable for QUHT.
As a corollary to [17, Lemma 1, Corollary 6], one could get a decision rule $\hat{M}_{m}$ such that
@@
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(\hat{M}_{m}))\right\}$ $\leq-\hat{D}(\rho,\sigma),$ (37)
where $\hat{D}(\rho\|\sigma)$ is a cousin of the relative entropy, defined as
$\hat{D}(\rho\|\sigma)=\lim_{s\to 0}\left\{\frac{-\ln\left(\operatorname{Tr}\left[\sigma^{\frac{1-s}{2s}}\rho\sigma^{\frac{1-s}{2s}}\right]^{s}\right)}{s}\right\}.$ (38)
We briefly show the result in Section A-J. However, [17, Lemma 3] showed that $\hat{D}(\rho\|\sigma)\leq D(\rho\|\sigma)$, and the relationship between $\hat{D}(\rho\|\sigma)$ and the trace norm is not known, which makes direct comparison difficult. Nevertheless, the quantum empirical distribution they define is likely worth consideration for establishing tighter bounds for QUHT.
@@
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta_{\rho,\sigma}(M_{m}^{*}))\right\}=-D(\rho\|\sigma),$ (39)
then applying Quantum Pinsker’s Inequality, which is tight up to a constant factor [^1] When $\rho,\sigma$ share an eigen basis, the inequality reduces to the classical version, which is known to be tight to the constant [43], we would have $\beta_{\rho,\sigma}(M^{*}_{m})\simeq e^{-mD(\rho,\sigma)}\leq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. However, in Theorem 6, we showed that there exists a decision rule $M_{m}$ such that $\beta_{\rho,\sigma}(M_{m})\simeq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. In other words, even if a decision rule achieving an error exponent of $D(\rho\|\sigma)$ were available, its implied performance in terms of trace distance would be upper-bounded by a term matching the exponent achieved using tomography.
Several other works have considered similar problems under different names in other areas of quantum information science [21] [22] [27] [28], often motivated by practical needs for verifying quantum devices or certifying quantum behavior in minimally structured settings.. By framing QUHT as a formal hypothesis testing problem, we have aimed to provide both conceptual clarity and a foundation for future work. While our focus has been on leveraging quantum state tomography to achieve provable error exponents, the broader framework of one- and two-sample QUHT may have applications analogous to those of classical universal hypothesis testing. We view the present results as a first step in this direction, and hope they stimulate further theoretical and applied developments.
@@ Appendix A Proofs
We also have the following quantum version of Pinsker’s inequality [42, Theorem 3.1]
@@
$D(\rho\|\sigma)\geq\frac{1}{2}\left\|\rho-\sigma\right\|_{1}^{2}.$
A-A Proof of Theorem 1
[…]
@@ A-A Proof of Theorem 1
$\hat{H}_{0}:$ $\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m$
$\hat{H}_{1}:$ $\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})<m.$
@@
$\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]=F(\rho,\sigma),$
@@
$\beta_{\rho,\sigma}(M_{m})$ $=\mathbb{P}\left[\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m\right]$
$=\left[\prod_{i=1}^{m}\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]\right]$
$=\left[F(\rho,\sigma)\right]^{m}.$
A-B Tightness of Symmetric Hypothesis Testing Error Bound

### v3

@@ III-C Qudit Independent Measurements
Unfortunately, extending the above result to higher-dimensional Hilbert spaces gives an error exponent that scales poorly with the dimension of the space. For example, any Hermitian operator on $b$ qubits ($d=2^{b}$) can be decomposed using $d^{2}-1$ components, from Pauli strings [40]. Then applying the same method would give
@@
However, for general $d$-dimensional qudit states, we can do better using independent measurements by applying the following result from QST [41]:
@@
$\mathbb{P}\left[\left\|\tau-\hat{\tau}\right\|_{1}\geq\delta\right]\leq d\exp\left\{-\frac{m\delta^{2}}{86r^{2}d}\right\}.$ (21)
Note that this is a special case of Guţă’s [41] more general result, where they take measurements that are $2$-designs [41, Definition 1], and bounding the operators rank by $d$. This yields a concentration result that we can use in the same manner as before.
@@
$\beta(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86d^{3}}\right\},$ (23)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\right)$ term in the exponent.
This also yields a result for the two-sample QUHT setting.
@@
$\beta(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344d^{3}}\right\},$ (25)
where $\lesssim$ is hiding an additive $O(k^{1/2})$ term in the exponent.
@@ III-D Entangled Measurements
While independent measurements are easier to implement, it has been shown that entangled measurements achieve a sample complexity that is strictly better than that achievable by independent measurements [42].
@@
$\beta(M_{m})$ $\lesssim\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\},$ (29)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\right)$ term in the exponent.
We can also use this to get results for two-sample QUHT. In particular, using entangled measurements, we can achieve the following:
@@
$\beta(M_{m})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\},$ (31)
where $\lesssim$ hides an additive $O(k^{1/2})$ term in the exponent.
@@ IV Discussion
Recall that in the classical universal setting, Hoeffding was able to characterize the error exponent in terms of the KLD between the nominal and true distributions. In this work, we have established positive error exponents not in terms of the KLD, but the trace distance between the nominal and true distributions. Specifically, we showed various results of the following form:
$\lim_{n\to\infty}\left\{\frac{1}{m}\ln(\beta(M_{m}))\right\}=-\frac{\left\|\rho-\sigma\right\|_{1}^{2}}{c},$ (32)
@@
Comparing with the classical results for the universal hypothesis testing problem, which achieves an error exponent in the form of the relative entropy between $\rho$ and $\sigma$, this result appears somewhat lacking. For example, we have quantum Pinsker’s inequality [43, Theorem 3.1], which establishes that the squared trace distance forms a lower bound for the relative entropy. Since Hoeffding establishes his result in the classical setting using a large deviation type bound, a first attempt at a quantum generalization might utilize quantum variations of the classical Sanov’s theorem [6, Theorem 11.4.1]. Unfortunately, the results in [11, 13] are insufficient for recovering a bound in the QUHT setting, as they require knowledge of the alternative hypothesis (the true state $\sigma$ in this work).
As a corollary to [14, Lemma 1, Corollary 6], one could get a decision rule $\hat{M}_{m}$ such that
@@
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta(\hat{M}_{m}))\right\}$ $\leq-\hat{D}(\rho,\sigma),$ (34)
where $\hat{D}(\rho\|\sigma)$ is defined as
$\hat{D}(\rho\|\sigma)=\lim_{s\to 0}\left\{\frac{-\ln\left(\operatorname{Tr}\left[\left(\sigma^{\frac{1-s}{2s}}\rho\sigma^{\frac{1-s}{2s}}\right)^{s}\right]\right)}{s}\right\}.$ (35)
We briefly show the result in Section˜A-F. However, [14, Lemma 3] showed that $\hat{D}(\rho\|\sigma)\leq D(\rho\|\sigma)$, and tight relationships between $\hat{D}(\rho\|\sigma)$ and the trace norm are not known, which makes direct comparison difficult. Nevertheless, the quantum empirical distribution they define is likely worth considering for establishing tighter bounds for QUHT.
@@
$\lim_{m\to\infty}\left\{\frac{1}{m}\ln(\beta(M_{m}^{*}))\right\}=-D(\rho\|\sigma),$ (36)
then applying Quantum Pinsker’s Inequality, which is tight up to the constant  [^1] When $\rho,\sigma$ commute, the inequality reduces to the classical version, which is known to be tight to the constant [44], we would have $\beta(M^{*}_{m})\simeq e^{-mD(\rho,\sigma)}\leq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. However, in Theorem˜6, we showed that there exists a decision rule $M_{m}$ such that $\beta(M_{m})\simeq e^{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}}$. In other words, even if a decision rule achieving an error exponent of $D(\rho\|\sigma)$ were available, its implied performance in terms of trace distance would be upper-bounded by a term matching the exponent achieved using tomography. We corroborate this in Section˜A-G where we show that any result achieving a constant type I error rate, and a type II error exponent of $C\left\|\rho-\sigma\right\|_{1}^{2}$ with $C>\frac{1}{2}$ would violate the converse result of classical Stein’s lemma.
Several other works have considered similar problems under different names in other areas of quantum information science [21] [22] [27] [28], often motivated by practical needs for verifying quantum devices or certifying quantum behavior in minimally structured settings. By framing QUHT as a formal hypothesis testing problem, we hope that it will provide both conceptual clarity and a foundation for future work. While our focus has been on leveraging QST to achieve provable error exponents, the broader framework of one- and two-sample QUHT may have applications analogous to those of classical universal hypothesis testing.
@@ Appendix A Proofs
We also have the following quantum version of Pinsker’s inequality [43, Theorem 3.1]
@@
$D(\rho\|\sigma)\geq\frac{1}{2}\left\|\rho-\sigma\right\|_{1}^{2}.$ (38)
A-A One-Sample Generic Result
Here we will prove a generic result which shows that any concentration result for quantum state tomography of a particular form gives us a test achieving a certain type two error exponent in the QUHT setting.
Lemma 6.
Given $\rho$ the density operator of $\sigma$ under $H_{0}$, suppose there is a method of constructing $\hat{\sigma}_{m}$ from $\sigma^{\otimes m}$ such that a concentration result of the following form holds: $\forall t>0$, and for a constant $C$ and a function $g(m)=\operatorname{poly}(m)$
$\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\geq t\right]\leq g(m)\exp\left\{-mCt^{2}\right\}$ (39)
Then using the test
$\hat{H}_{0}$ $:\left\|\hat{\sigma}_{m}-\rho\right\|_{1}\leq c_{m}$ (40)
$\hat{H}_{1}$ $:\left\|\hat{\sigma}_{m}-\rho\right\|_{1}>c_{m}$ (41)
Given $\alpha_{m}>0$, if we took $c_{m}=O\left(\frac{\ln^{1/2}(1/\alpha)}{\sqrt{m}}\right)$, we would get
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]\leq\alpha$ (42)
And for $m$ large enough,
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]\leq g(m)\exp\left\{-m\ C\left\|\rho-\sigma\right\|_{1}^{2}+O\left(m^{1/2}\right)\right\}$ (43)
Proof.
The type I error can be written in terms of the probability of having $\hat{\sigma}_{m}$ far from $\sigma$, which we can bound directly using the concentration result. Specifically, we have
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ $=\mathbb{P}\big[\left\|\hat{\sigma}_{m}-\rho\right\|_{1}>c_{m}\ \big|\ \sigma=\rho\big]$ (44)
$=\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}>c_{m}\right]$ (45)
$\leq g(m)\exp\left\{-m\ C\ c_{m}^{2}\right\}$ (46)
Then taking
$c_{m}$ $=\left(\frac{\ln(g(m)/\alpha)}{Cm}\right)^{1/2}$ (47)
$=O\left(\frac{\ln(g(m))\cdot\ln^{1/2}(1/\alpha)}{m^{1/2}}\right)$ (48)
We would get $\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]\leq\alpha$.
For the type II error, we have
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ $=\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\rho\right\|_{1}\leq c_{m}\right]$ (49)
$\leq\mathbb{P}\left[\left\|\sigma-\rho\right\|_{1}-\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\leq c_{m}\right]$ (50)
$=\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\geq\left\|\sigma-\rho\right\|_{1}-c_{m}\right]$ (51)
And since $\rho\neq\sigma$ and $\sigma,\rho$ fixed in this setting, and since $c_{m}\to 0$ as $m\to\infty$, then for $m$ large enough $c_{m}<\left\|\rho-\sigma\right\|_{1}$. When this condition is satisfied, we have
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]\leq g(m)\exp\left\{-Cm\left(\left\|\rho-\sigma\right\|_{1}-c_{m}\right)^{2}\right\}$ (52)
$\leq g(m)\exp\left\{-C\ m\left(\left\|\rho-\sigma\right\|_{1}^{2}-2c_{m}\right)\right\}$ (53)
$=\exp\left\{-m\ C\left\|\rho-\sigma\right\|_{1}^{2}+\frac{2m^{1/2}\ln^{1/2}(1/\alpha)}{C^{1/2}}+\ln(g(m))\right\}$ (54)
Note that Theorem˜2 and Theorem˜4 are direct corollaries of this result using Lemma˜1 and Lemma˜2 respectively. And as corollary of Lemma˜3, we have the following concentration result in terms of the trace norm:
Corollary 1.
Using Lemma˜3, we can construct $\hat{\sigma}_{m}$ using $\sigma^{\otimes m}$ such that
$\mathbb{P}\left[\left\|\sigma-\hat{\sigma}_{m}\right\|_{1}\geq t\right]\leq(m+1)^{3d^{2}}\exp\left\{-\frac{mt^{2}}{2}\right\}$ (55)
Proof.
$\mathbb{P}\left[F(\sigma,\hat{\sigma}_{m})\leq 1-\delta\right]$ $\geq\mathbb{P}\left[1-\frac{1}{4}\left\|\sigma-\hat{\sigma}_{m}\right\|_{1}^{2}\leq 1-\delta\right]$ (56)
$=\mathbb{P}\left[\left\|\sigma-\hat{\sigma}_{m}\right\|_{1}\geq 2\delta^{1/2}\right]$ (57)
$\implies\mathbb{P}\left[\left\|\sigma-\hat{\sigma}_{m}\right\|_{1}\geq t\right]$ $\leq(m+1)^{3d^{2}}\exp\left\{-\frac{mt^{2}}{2}\right\}$ (58)
And Corollary˜1 combined with Lemma˜6 gives us Theorem˜6.
A-B Two-Sample Generic Result
And in a similar manner, we can prove that the same type of generic concentration result yields a test with a generic bound on the type II error.
Lemma 7.
Given the same general concentration result as in Equation˜39 for the two sample setting with $k=\min\{m,n\}$, we could use the decision region
$\hat{H}_{0}$ $:\left\|\hat{\sigma}_{m}-\hat{\rho}_{n}\right\|_{1}\leq c_{k}$ (60)
$\hat{H}_{1}$ $:\left\|\hat{\sigma}_{m}-\hat{\rho}_{n}\right\|_{1}>c_{k}$ (61)
We assume that $\limsup\frac{n}{m}=a<+\infty$, i.e. that $n$ and $m$ go to $\infty$ with bounded ratio.
And given $\alpha>0$, taking $c_{k}=O\left(\frac{\ln((g(n)+g(m))/\alpha)}{k^{1/2}}\right)$ would yield
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]\leq\alpha$ (62)
And
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]\leq(g(n)+g(m))\exp\left\{-k\frac{C\left\|\rho-\sigma\right\|_{1}^{2}}{4}+O\left(k^{1/2}\right)\right\}$ (63)
Proof.
$\mathbb{P}\left[\hat{H}_{1}|H_{0}\right]$ (64)
$=\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\hat{\sigma}_{n}\right\|_{1}>c_{k}\right]$ (65)
$\leq\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}+\left\|\hat{\sigma}_{n}-\sigma\right\|_{1}>c_{k}\right]$ (66)
$\leq\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}>\frac{c_{k}}{2}\right]+\mathbb{P}\left[\left\|\hat{\sigma}_{n}-\sigma\right\|_{1}>\frac{c_{k}}{2}\right]$ (67)
$\leq g(m)\exp\left\{-m\ C\frac{c_{k}^{2}}{4}\right\}+g(n)\exp\left\{-mC\frac{c_{k}^{2}}{4}\right\}$ (68)
$\leq\left(g(n)+g(m)\right)\exp\left\{-kC\frac{c_{k}^{2}}{4}\right\}$ (69)
where $k=\min(m,n)$.
Then if we have
$c_{k}$ $=\sqrt{\frac{4\ln\left(\frac{g(n)+g(m)}{\alpha}\right)}{Ck}}$ (70)
$=O\left(\frac{\ln^{1/2}(1/\alpha)}{k^{1/2}}\right)$ (71)
Where we can ignore the $\ln(g(n)+g(m))$ terms so long as as $\limsup\frac{n}{m}=a<+\infty$, so that $\ln(g(n)+g(m))/k\to 0$. Then $c_{k}\to 0$ as $k\to\infty$
Now, targeting the type II error, we assume that $\rho$ and $\sigma$ are fixed, and $\rho\neq\sigma$.
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ (72)
$=\mathbb{P}\left[\left\|\hat{\rho}_{n}-\hat{\sigma}_{m}\right\|_{1}\leq c_{k}\right]$ (73)
$\leq\mathbb{P}\left[\left\|\rho-\sigma\right\|_{1}-\left\|\hat{\rho}_{n}-\rho\right\|_{1}-\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\leq c_{k}\right]$ (74)
$=\mathbb{P}\left[\left\|\hat{\rho}_{n}-\rho\right\|_{1}+\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\geq\left\|\rho-\sigma\right\|_{1}-c_{k}\right]$ (75)
$\leq\mathbb{P}\left[\left\|\hat{\rho}_{n}-\rho\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{k}}{2}\right]$ (76)
$\ \ +\mathbb{P}\left[\left\|\hat{\sigma}_{m}-\sigma\right\|_{1}\geq\frac{\left\|\rho-\sigma\right\|_{1}-c_{k}}{2}\right]$ (77)
And since $\rho\neq\sigma$, for $k$ large enough we would have $\left\|\rho-\sigma\right\|_{1}-c_{k}>0$. Then using the concentration result, we get
$\mathbb{P}\left[\hat{H}_{0}|H_{1}\right]$ (78)
$\leq g(n)\exp\left\{-nC\frac{(\left\|\rho-\sigma\right\|_{1}-c_{k})^{2}}{4}\right\}$ (79)
$\ \ \ +g(m)\exp\left\{-mC\frac{(\left\|\rho-\sigma\right\|_{1}-c_{k})^{2}}{4}\right\}$ (80)
$\leq\left(g(n)+g(m)\right)\exp\left\{-kC\frac{\left\|\rho-\sigma\right\|_{1}^{2}-2c_{k}}{4}\right\}$ (81)
And using Lemma˜7, we have Theorem˜3 Theorem˜5 and Theorem˜7 as corollaries of using Lemma˜1 Lemma˜2 and Lemma˜3 respectively.
A-C Proof of Theorem˜1
Since $\rho$ is a pure state, there exists a unitary operator $U\in\mathcal{L}(\mathcal{H})$ such that $U\rho=\lvert 0\rangle\langle 0\rvert$. Similarly, we can apply $U^{\otimes m}$ to $\sigma^{\otimes m}$, such that without loss of generality we can consider $\rho$ to be the pure state $\lvert 0\rangle\langle 0\rvert$ (similarly, in the one sample case we can always assume without loss of generality that $\rho$ is diagonalizable in the computational basis).
@@ A-C Proof of Theorem˜1
$\begin{split}\hat{H}_{0}:&\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m\\ \hat{H}_{1}:&\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})<m.\end{split}$ (82)
@@
$\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]=F(\rho,\sigma),$ (83)
@@
$\beta(M_{m})$ $=\mathbb{P}\left[\sum_{i=1}^{m}M_{\lvert 0\rangle}(\sigma_{i})=m\right]$ (84)
$=\left[\prod_{i=1}^{m}\mathbb{P}\left[M_{\lvert 0\rangle}(\sigma_{i})=1\right]\right]$ (85)
$=\left[F(\rho,\sigma)\right]^{m}.$ (86)
A-D Tightness of Symmetric Hypothesis Testing Error Bound